OpenStudy (big_mensa961):

If the equation of a circle is (x + 5)2 + (y - 7)2 = 36, its radius is 6 16 36

OpenStudy (studygurl14):

Equation of a circle: \(\large x^2+y^2=r^2\) where r = radius

OpenStudy (studygurl14):

EDIT: \(\large (x-h)^2+(x-k)^2=r^2\)

OpenStudy (dinamix):

r^2=36 so r= ? @Big_Mensa961